OpenStudy (anonymous):

Some psychologists believe that a "genius" should be defined as anyone having an IQ over 140. If IQ scores are normally distributed with a mean of 100 and a standard deviation of 15, and if the population of the world is 6,575,000,000, how many geniuses are there in the world today?

OpenStudy (amistre64):

so your wanting to find the % of n ..... so find the probability of greater than 140

OpenStudy (amistre64):

z = (x-mean)/sd; and table it out or use the normalCDF function on a ti83

OpenStudy (anonymous):

I tabled it out and got .9961

OpenStudy (amistre64):

id have to check the table ... :) 40/15 = 2.66666 so thats between 95 and 99.something

OpenStudy (anonymous):

ok so what's the next step now

OpenStudy (amistre64):

.9961 are below, so 1-.9961 are above

OpenStudy (amistre64):

so at 140, left tail is .9961, right tail has to be .0039

OpenStudy (amistre64):

n*.0039 is the amount of n above 140

OpenStudy (anonymous):

so then we multiply .0039 by the total population to find the number of geniuses?

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

if the population is n, and we determined the percentage ... then n*% is the solution

OpenStudy (anonymous):

So i got 25,642,500 when I put it in my calculator but that's not any of the answer choices

OpenStudy (amistre64):

thats correct, but i got no idea what your options are

OpenStudy (anonymous):

one of the answers is close but it's off by like a million

OpenStudy (anonymous):

A. 24,985,000 B. 536,000,000 C. 123,000 D. 115,000 E. 6,575,000

OpenStudy (amistre64):

id go with the 24.9 then since its the closest one

OpenStudy (anonymous):

alright. thanks for your help!

OpenStudy (amistre64):

yep dblckd ... z = (140-100)/15 = 2.6666 since its a right tail, your table would give the complement as -2.6666 as well

OpenStudy (anonymous):

which comes out as .0039 got it

OpenStudy (amistre64):

http://ucce.ucdavis.edu/files/repository/calag/tab4605p15.jpg even a right tailed table has us at .0039 .... so all signs point to something close to 25m